the length of a rectangle is 100ft. its area is twice the perimeter. How wide is the rectangle?
Answer by lwsshak3(11628)

the length of a rectangle is 100ft. its area is twice the perimeter. How wide is the rectangle?
***
let x=width of the rectangle
perimeter=2*length+2*width=200+2x
Area=length*width=100x
area=2*perimeter
100x=2(200+2x)
100x=400+4x
96x=400
x=400/96≈4.17
How wide is the rectangle? 4.17 ft